OpenStudy (anonymous):

how to find the equation of a tangent line to the graph x^3 + y^3=6xy

OpenStudy (primeralph):

First find the derivative with respect to x.

OpenStudy (anonymous):

and that is \[3x ^{2}+3y ^{2}=1\]

OpenStudy (anonymous):

or do i differentiate expricitly?

OpenStudy (primeralph):

You have to do it implicitly.

OpenStudy (anonymous):

but i still didnt do it right?

OpenStudy (primeralph):

No you didn't yet.

OpenStudy (anonymous):

\[\left(\begin{matrix}d \\ dx\end{matrix}\right)\left[ x ^{3}+y ^{3}\right]=\left(\begin{matrix}d \\ dx\end{matrix}\right)\left[ 6xy \right]\]

OpenStudy (anonymous):

so \[\left(\begin{matrix}dy \\ dx\end{matrix}\right)= \frac{ 3x ^{2} }{ -2x ^{2}}\]

hartnn (hartnn):

do you know about chain rule ? d/dx (y^3) = .... ? (requires chain rule)

OpenStudy (anonymous):

ugh okay so then do i do the chain rule from the very beggining??

hartnn (hartnn):

d/dx(x^3) does not require chain rule d/dx(6xy) requires chain and product rule

OpenStudy (anonymous):

well what if i just moved the 6x?

hartnn (hartnn):

moved where ? keeping it as 6xy and applying chain+product rule is the easiest thing you can do here

hartnn (hartnn):

what do u get for d/dx (y^3) =... ?

OpenStudy (anonymous):

3y

OpenStudy (anonymous):

well \[3y ^{2}\]

hartnn (hartnn):

d/dy(y^3) = 3y^2 but here we are trying to find d/dx(y^3) notice the difference ?

OpenStudy (anonymous):

so what am i doing then is it just 3y dy/dx?

hartnn (hartnn):

if you mean \(3y^2 \dfrac{dy}{dx}\) , then yes! what about 6xy ? know product rule?

OpenStudy (anonymous):

so what is f(x) and what is g(x) if there are three different variables instead of two? or is f(x) 6x and g(x) y

hartnn (hartnn):

6 is kust a constant, you can throw it out of the differentiation... d/dx (6xy) = 6 d/dx(xy) try to find d/dx(xy)

OpenStudy (anonymous):

isnt it just 1?

OpenStudy (anonymous):

wait no never mind

OpenStudy (anonymous):

x+y

hartnn (hartnn):

product rule, \((fg)' = f'g+fg'\) \((xy)' = x'y+xy' \) whats x' = ... ?

OpenStudy (anonymous):

1

hartnn (hartnn):

yes, so, 6 d/dx(xy) is just 6 (y+ x dy/dx) got this ?

OpenStudy (anonymous):

but isnt the rule f(x)g'(x)+g(x)f'(x)?? so then it would be x(1)+y(1) therefore x+y??

hartnn (hartnn):

\(\large y' = \dfrac{dy}{dx}\) because we are differentiating with respect to y

hartnn (hartnn):

i mean we are differentiating with respect to x ***

OpenStudy (anonymous):

so then the dy/dx goes next to the y right?

hartnn (hartnn):

yes, so after your differentiation implicitly, you get \(\Large 3x^2+3y^2 \dfrac{dy}{dx} = 6 (x\dfrac{dy}{dx}+y)\) any doubts till here ?

OpenStudy (anonymous):

i have that

OpenStudy (anonymous):

so then i need to isolate the dy/dx on one side and factor it out?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

how can i do that? do i divide by 6 first?

hartnn (hartnn):

dividing by 6 will give you fractions... lets do , divide by 3 on both sides.

hartnn (hartnn):

then you can move y^2 dy/dx from left to right

OpenStudy (anonymous):

why 3??

hartnn (hartnn):

to avoid fractions... if you divide by 6, on left side you get first term as (1/2) x^2

OpenStudy (anonymous):

why dont i destribute the 6?

hartnn (hartnn):

you surely can do that... and in later steps, you will realize that you need to divide 3 on both sides... if you do it first or later, it will make no difference...

OpenStudy (anonymous):

okay so after the three division?

hartnn (hartnn):

\(\Large x^2+y^2 \dfrac{dy}{dx} = 2 (x\dfrac{dy}{dx}+y)\) distribute the 2

OpenStudy (anonymous):

then subtract 2y?

hartnn (hartnn):

yes what do u get ?

OpenStudy (anonymous):

\[x ^{2}+y ^{2}\left(\begin{matrix}dy \\ dx\end{matrix}\right)-2y=2x \left(\begin{matrix}dy \\ dx\end{matrix}\right)\]

hartnn (hartnn):

now to combine terms with dy/dx you can subtract y^2 dy/dx from both sides, try it

OpenStudy (anonymous):

\[x ^{2}-2y=\left(\begin{matrix}dy \\ dx\end{matrix}\right)\left( 2xy ^{2} \right)\]

OpenStudy (anonymous):

and then divide by 2xy^2

hartnn (hartnn):

nooooooooooooo

OpenStudy (anonymous):

why not

hartnn (hartnn):

when you subtract, it'll be \(\Large 2x-y^2\) on right side

hartnn (hartnn):

and not 2xy^2

OpenStudy (anonymous):

wait why am i subtracting y^2?

OpenStudy (anonymous):

isnt it multiplied by 2x

hartnn (hartnn):

to combine terms with dy/dx together \(x ^{2}+y ^{2}\left(\begin{matrix}dy \\ dx\end{matrix}\right)-2y=2x \left(\begin{matrix}dy \\ dx\end{matrix}\right) \\ x ^{2}+y ^{2}\left(\begin{matrix}dy \\ dx\end{matrix}\right)-y ^{2}\left(\begin{matrix}dy \\ dx\end{matrix}\right)-2y=2x \left(\begin{matrix}dy \\ dx\end{matrix}\right) -y ^{2}\left(\begin{matrix}dy \\ dx\end{matrix}\right)\\ x ^{2}-2y=(2x-y^2) \left(\begin{matrix}dy \\ dx\end{matrix}\right)\) see if u get this ?

hartnn (hartnn):

in the last step , i factored out dy.dx so what remained was 2x and -y^2 so, (2x-y^2)

OpenStudy (anonymous):

alright i see what you did... and noooow you divide by 2x-y^2?

hartnn (hartnn):

thats absolutely correct! and you will get your required dy/dx :)

OpenStudy (anonymous):

and that is the slope to my tangent line?

OpenStudy (anonymous):

are you sure? because the slope should = -1

hartnn (hartnn):

at which point do you need the slope ?

OpenStudy (anonymous):

(3,3)

hartnn (hartnn):

slope at 3,3 = dy/dx at 3,3 so put x= 3 and y=3 in the derivative dy/dx which you got

OpenStudy (anonymous):

i think it would have been easier if we distributed the 6x first because then i could have moved dy/dx 6x to the left and from the left taken 3x and put that on the right

OpenStudy (anonymous):

then i would have all the dy/dx on one side and factor it out and divide by that on the right and simplify into the -1/2

OpenStudy (anonymous):

well -1

hartnn (hartnn):

whatever you do, if you get dy/dx = (x^2-2y)/ (2x-y^2) then you'll be correct, else incorrect.

hartnn (hartnn):

and good, you got the slope as m=-1 as required :)

OpenStudy (anonymous):

i wish i could show you the work.... i dont know why i didnt see it sooner

OpenStudy (anonymous):

anyways thanks for your help and being patient and stuff.... sorry for taking up so much of your time

hartnn (hartnn):

hey, no problem :) always happy to help :) welcome ^_^

OpenStudy (anonymous):

i ended with dy/dx= 6y-3x/ 3y-6x and it simplifies down to negative one

hartnn (hartnn):

oh, haha...now you know the correct dy/dx right ? :) ask if u have any more doubts on this ?

OpenStudy (anonymous):

i know the correct answer and all but i have a test on derivatives, rates of change, and related rates and i still dont understand it all and the test is tomorrow so im not very confident about it

hartnn (hartnn):

don't worry much...practice 2 or 3 questions on each topic, and you'll be good to go :)

OpenStudy (anonymous):

but the test is tomorrow and im already really stressed because im baking a pie

OpenStudy (anonymous):

for pi day tomorrow

OpenStudy (primeralph):

A real math person ^^^.

OpenStudy (anonymous):

then why do i suck at math so much

OpenStudy (anonymous):

i remember i used to like math and now im in calculus ksdjfihafiunoisdf

OpenStudy (primeralph):

You just have to find how to understand the basics in a way that suits you; then you'll be golden.